NEWS

These are the news starting 01 Jan 2014
Click here for the news of the previous years.

31 Dec 14 These are the last news for this year 2014!
1. Overplot of total ozone column measurements done at Uccle and meteoLCD now on the Dobson page (here). This gives a first impression of the similarity of both measurement series.
2. Next 7 days graphs will be corrupted due to the usual Gnuplot problem. This proble is sel-healing in 7 days:-)))
3. meteoLCD wishes all its friends and visitors a very Happy New Year.
02 Dec 14 Upload of November data (nov14.dat file) into the data archive.
28 Nov 14 There is a possible malfunction of the main pump which sucks in atmospheric air for distribution to the ozone and CO2 sensors. Starting 26Nov14, the daily variations of these 2 gases seem too small.
01 Nov 14 1. Maintenance on UVB and UVA sensors (cleaned contacts, changed dessicant).
2. Upload of October data (oct14.dat file) into the data archive.
3. The logger has been emptied and restarted. The existing raw data file all.dat has been saved (as all_26Aug13to31Oct14.dat) into the data archive. A new all.dat file ( pointed at by the link "all data since last logger restart") will be generated at midnight.
4. The spurious peaks in wind speed (ch.61 and ch. 4) should probably be ignored.
31 Oct 14 Recalibration of CO2 sensor and re-check of O3 sensor. The UVA sensor is out of work for about 1 day. The inlet filter for all gas sensors was very dirty; has been cleaned.
24 Oct 14 Maintenance by T.S.E. 
The AC31M NO/NO2 sensor which is out of order since more than a month has been removed.. This will possibly be the end of the meteoLCD NOx measurements after more than 16 years of operation, as we can not afford a new sensor and have to concentrate on our most important gas measurements (which clearly are the CO2 measurements).
23 Oct 14 A new paper by F.Massen and R.Tholl showing how to calibrate the new Pt100 sensor of the Hygrosmart module, which replaces the defunct thermistor previously used for measuring ambient air temperature. The data file will hold in Air Temperature channel 2 the raw Ohm readings of the Pt100, and the calibration function given in the paper should be used to transform these Ohm readings into degrees Celsius.
01 Oct 14 Upload of September data (sep14.dat) into data archive.
20 Sep 14 Maintenance and calibration work by Raoul.
03 Sep 14 Upload of August 2014 data into data archive.
Please read metadata at the start of the file with care!
25 Aug 14 1. Cleaning of connectors of the UV sensors. The Peltier element of the UVB now is ok.
2. The NO/NO2 sensor AC31M (Environnement SA) has been shut down and will remain off until the next maintenance by T.S.E. The malfunction is due to a non-working Peltier element.
24 Aug 14 1. The NO2/NO sensor is out of order (the Peltier element in the reaction chamber seems to be dead). No real hopes to find a quick remedy!
2. The internal temperature stabilization of the UVB sensor also is down, also probably due to a Peltier failure.This can possibly be brought back to life tomorrow.
21 Aug 14 1. Calibration and maintenance.
2. Mounted with Raoul Tholl a support for an external stand-alone ONSET temperature logger, to be used for a calibration of the Pt100 sensor.
12 Aug 14 Trying to use the combined PT100/RH Hydrosmart sensor from Michell that Raoul mounted the 25th July. This sensor gave bad air temperature readings as the data logger expected a thermistor (and not a PT100). I changed the calibration of CH2 today to measure resistance, so the numbers appearing under the label will be OHM's. Note that 100 Ohm correspond to 0 C and that 30 C to 111.672 Ohm. The resistance rises approx. 0.385 Ohm above 100 Ohm per C. We will leave this for a time and than try a calibration factor using the Drybulb sensor and the temperature sensor of the backup Davis VantagePro station. Changing  the configuration of the datalogger requires a data dumb and restart from fresh.
08 Aug 14 The main page with the near-live data had some dead links. These have been corrected, together with some minor changes.
01 Aug 14 Uploaded jul14.dat to data archive. Please read comments at the start of the file carefully!
31 Jul 14 1. Dobson data uploaded; measurements since 15/07 are made by Mike Zimmer
2. Live plots of Windchill corrected to use the replacement DryBT sensor for air temperature.
25 Jul 14 This was sort of a black Friday. A computer malfunction stopped all transmissions from about 12:00 to late after-noon. Thx to Raoul's efforts everything is working again. No data are lost, as everything is stored in the separate datalogger before being uploaded to the webserver.
24 Jul 14 1. Starting 10:30 UTC the drybulb_T sensor is used as a replacement for the non-working usual air_temperatur sensor. All the graphs involving air temperature are built on this replacement sensor readings (today as well the 7days graphs). A short calibration check between the AirT and Drybulb sensors shows this. I suggest that we keep to the linear calibration curve which means the readings of both sensors are essentially (statistically) the same.
2. The new version of L2K hang at midnight. Restarted the program... seems ok now. Will keep an eye on this.
23 Jul 14 1. Raoul Tholl replaced the T/H sensor by a spare module, which after a short time gives impossible temperature readings (rel. humidity may be ok). If this problem remains, the readings from the WBGT sensor will be used as a stop-gap.
2. New version of the Lightning2000 (L2K) software installed.
17 Jul 14 The relative humidity sensor is dead. This means that the moist enthalpy graph should be ignored, as the humidity is one of the parameters used to calculate moist enthalpy. A new sensor will probably not be available before mid-August. Sorry!
10 Jul 14 Calibration of the CO2 and NO/NO2 sensors.
07 Jul 14 Lightning computer running ok now (06:30 UTC). The computer hang during a browser update (which should not have been done!)
06 Jul 14 There is a problem with the lightning computer...try to solve this asap.
01 Jul 14 Uploaded jun14.dat file into data archive.
19 Jun 14 1. Continued the imaging research: finally I got the best results using the freely available PARAGON Back & Recovery 2014 (free edition), which I put on a bootable CD. Imaging in raw data worked fine.
2. Tried that Paragon software with other machines: XPPro (bureau): OK, Win2k (radioactivity): OK.
18 Jun 14 1. The lightning sensor behaves badly: far away storms are shown too close to the Luxembourg border. To get to the problem, I first dismounted the antenna and cleaned the RJ45 connector. We will wait for the next storm to make a further decision i.e. roll back the software Lightning2000 to an older version.
2. I had many problems with imaging the main Linux machine (which does most of the stuff). Tried several solutions and a new larger ext. USB disk, with bad results.
Expect some disturbances in the graphs!
13 Jun 14 The WBGT kludge did not hold very long: probably the fittings came loose the 8th June and the wet bulb pot dried out. I replaced today at 13:40 the whole upper section of the tubing, adding a second silicon tube as a protection against bird pickings. Lets hope that this repair will hold better than the previous one :-))
07 Jun 14 During a routine check I saw that the WETBULB sensor (CH.21) was dry. The plastic tube bringing de-mineralized water to the wetbulb sensor was broken, so that pumped water did not reach the pot holding the wet wick. This malfunction started the 04Jun14 (probably during the morning). I made a quick kludgy repair (at 08:30 UTC, today), but the tube which has become very brittle must be replaced asap.
01 Jun 14 Uploaded May data may14.dat into data archive.
22 May 14 Maintenance and recalibration work by Raoul Tholl.
05 May
14
Uploaded April 2014 data file (apr14.dat) into data archive.
26 Apr 14 1. Upload of Microtops II data file (26Apr14.mt5) into data archive. This file holds all the original raw measurement data of the TOC, AOT and atmospheric water column from 15 June 2013 to 26 April 2014.
2. New comment and results on radon washout in the BLOG.
24 Apr 14 1. The meteo.lcd.lu website has migrated to a new Restena web server. Many thx to the Restena staff for the good work!  (24 April 2014, 08:10 UTC). 
2. Maintenance and recalibration.
03 Apr 14 Uploaded March data mar14.dat into data archive.
02 Apr 14 1. I salvaged the pump from the disused API O3 sensor and replaced the noisy one of the CO2 sensor with this more smoothly running piece. Pumps are really a big big headache here at meteoLCD!
2. The NO/NO2 sensors has been recalibrated.
3. The WBGT sensor is back to action... may be it is a bit early for measuring heat stress, but I am a firm believer in the IPCC consensus that we feel global warming everywhere (tongue in cheek!), so be prepared!
29 Mar
14
1. The API O3 sensor shows its infamous "communication with relay board" error again, and gives absilute faulty readings. It has been replace at 08:00 with the O341M from Environnement SA. As this sensor has a different calibration factor, the logger had to be stopped, emptied of stored data and re-launched.
2. The NO/NOx sensor AC31M gives very dubious readings. NO could be ok, but NO2 is absolutely off (impossible high readings)... It could be that the photomultiplier which was shortly exposed to dim light will take some time to come back to normal operation.
28 Mar 14 1. Maintenance by T.S.E.
2. The AC31M has a new chopper motor, and is online again.
3. The  original O3 sensor from API Teledyne is back, but only time will tell if it is working ok.
08 Mar 14 Uploaded February data (feb14.dat) into data archive.
03 Mar 14 Calibration and maintenance work together with R. Tholl. The NOx sensor will be off for a longer time...
19 Feb 14 The NOX sensor is out of order (the chopper motor stopped functioning about a week ago).The relevant icons in the live data web page will be marked accordingly.
07 Feb 14 The update of the trends page to include the 2013 data is done. This is an interesting but time consuming job.
01 Feb 14 Upload of January 2014 data into the data archive.
31 Jan 14 1. Maintenance by T.S.E
2. As the API O3 sensor was bad since a couple of weeks, it is replaced by the old O341M from Ennvironnement SA. This sensor which was dormant for over 6 years seems to work reasonably well. The new sensor started at 12:30 UTC (first readings appear at 15:30 UTC, as logger was restarted at 15:00)
16 Jun 14 I had to replace the batteries from two UPS; this was not easy, as the old batteries had developed a bulge and could only be removed by opening the cabinet and using brute force and some metal bending. Interruption from 08:00 to 09:15..
12 Jan 14 I continue updating the "TRENDS" page. This is quite a lot of work, as it has to be done by hand with a lot of secondary sources to check. The main conclusion form the air temperature series is that there definitively is no warming here since 2002, and that yearly mean temperatures and winters become colder.
11 Jan 14 Maintenance and recalibration. The ground ozone sensor displays again an annoying error; I send quite a time looking for some bad contacts, but could not find any. Be careful with these ozone readings.
09 Jan 14 The rain sensor was blocked again by bird droppings; I do not understand how they make it, as the rim of the sensor is really sharp and should hinder them to siz down and drop their sh** into the sensor. Maybe they do it during flight, diving down like a Stuka :-))).
The blockage probably started the 27 Dec. 2013 and finished today at 08:30.
I used the remaining water in the sensor to estimate that the total precipitation from 27/12/13 to 08/01/14 was 35mm. The DAVIS rain sensor (backup station) shows 48.3mm; the difference comes mostly from a rain shower the 28/12/2013, which was missed by the main rain sensor. The Dec2013 (27 to 31) and Jan2014 (01 to 08) archive data will be corrected by using the DAVIS precipitation readings.
01 Jan 14 1. The usual new year problem with the 7 days graphs.. as the logger file does not contain the year, Gnuplot gets confused the 1rst week of every new year. This will self correct the 7th January.

Here are the channels from left to right in the data file that should be discarded:
 
channel Label Comment channel to use instead
7 UVA old Delta_T sensor ch. 10
8 CO2 old MIR9000 sensor ch.11

The main reason for keeping the data files unchanged is that a lot of users pick data from these files and rely on a stable format.

2. Uploaded dec13.dat file to data archive and 2013.zip, which is simply the concatenated file holding all monthly .dat files
 

 


Back Home